OpenStudy (anonymous):

Solve 0.7^x = 22. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

How do i do this?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

take log both sides

ganeshie8 (ganeshie8):

\(\large 0.7^x = 22\) \(\large \log 0.7^x = \log 22\) \(\large x \log 0.7 = \log 22\) \(\large x =\frac{ \log 22}{\log 0.7}\)

ganeshie8 (ganeshie8):

use ur calculator

OpenStudy (anonymous):

ahhyes i got it -8.6..... thanks

ganeshie8 (ganeshie8):

-8.66627 rounded to hundredth becomes -8.67